OpenStudy (anonymous):

Solve the proportion. x+1/3 = x/2

OpenStudy (anonymous):

Multiply both sides with 6 6(x+1/3)=6(x/2) 6x+2=3x 6x+2-3x=3x-3x 3x+2=0 3x+2-2=0-2 3x=-2 x=-2/3